OpenStudy (anonymous):

Solve, using linear combination. 3x – y = 4 2x + y = 6 (0, –4) (–1, –7) (2, 2) (3, 0)

OpenStudy (phantomcrow):

Solve both equations for y and set the results equal to each other to solve for x. When you get what x is, plug the answer back into one of the equations in terms of y and you will get y.

OpenStudy (anonymous):

@ethan726

pooja195 (pooja195):

Still need help?

OpenStudy (anonymous):

yes

OpenStudy (kendricklamar2014):

\[5x = 10 \] \[5 \times ? = 10\] 4 + y = 6 \[4 + ? = 6\] Can you answer the ? @WildQueen

OpenStudy (anonymous):

I DONT GET IT

OpenStudy (kendricklamar2014):

5 multiplied by what number equals 10 @WildQueen

OpenStudy (anonymous):

ok is it c

OpenStudy (kendricklamar2014):

Yes the answer is C because: \[5 \times 2 = 10\] and \[4 + 2 = 6\] which give you the coordinates: \[x = 2 ,~ y = 2\]

OpenStudy (kendricklamar2014):

Did you understand?

OpenStudy (anonymous):

yea